# Defending Against Physical Adversarial Patch Attacks on Infrared Human Detection

### v2

@@ Abstract
Infrared detection is an emerging technique for safety-critical tasks owing to its remarkable anti-interference capability. However, recent studies have revealed that it is vulnerable to physically-realizable adversarial patches, posing risks in its real-world applications. To address this problem, we are the first to investigate defense strategies against adversarial patch attacks on infrared detection, especially human detection. We have devised a straightforward defense strategy, patch-based occlusion-aware detection (POD), which efficiently augments training samples with random patches and subsequently detects them. POD not only robustly detects people but also identifies adversarial patch locations. Surprisingly, while being extremely computationally efficient, POD easily generalizes to state-of-the-art adversarial patch attacks that are unseen during training. Furthermore, POD improves detection precision even in a clean (i.e., no-attack) situation due to the data augmentation effect. Evaluation demonstrated that POD is robust to adversarial patches of various shapes and sizes. The effectiveness of our baseline approach is shown to be a viable defense mechanism for real-world infrared human detection systems, paving the way for exploring future research directions.
@@ 3.2 Detecting Patch-based Occlusions
Additionally, since the primary objective is to detect people robustly, we employ weighted cross-entropy loss [phan2020resolving_wce] to prioritize the human class over the patch class:
@@ 4.4 Comparison between Adversarial Training Scheme
Furthermore, we considered a POD variant incorporating an adversarial training scheme that explicitly uses adversarial patches during training, inspired by Ad-YOLO [16]. We denote this adversarial variant as “Adv-POD.” Note that the original Ad-YOLO is inadequate against location-optimized adversarial patches, such as HCB and shape-loc ones, because Ad-YOLO places adversarial patches only at the center of a person during training. The procedure for Adv-POD comprises four steps: (1) a model is trained using a history of adversarial patches; (2) a new adversarial patch is generated every 15 epochs, starting from the $5_{th}$ epoch; (3) the adversarial patches are stored in the adversarial patch history for subsequent training; (4) the stored patches are randomly selected and attached to input image to train the detection model. We used the AdvPatch attack [9] described above for adversarial patch generation. We positioned the adversarial patches in the exact same way as for POD, i.e., randomly determining the sizes and locations of patches for fair comparison.
@@ 5 Results
In this section, we present experimental results demonstrating the effectiveness of POD against diverse adversarial patch attacks.
@@
Patch Type Patch Detect. Clean ${}^{\dagger}$oise Patch ${}^{\dagger}$dvPatch HCB Time (rel.)
@@
Table 1: Average precision (AP@0.5) for adversarial patches. Model names with “${}_{noDet}$” indicate models without the patch-detection module described in Sec. 3.2. Attacks marked with $\dagger$ are not physically realizable in the infrared scenario. Results marked with $\ast$ are the results for seen attacks: the adversarial patch was encountered during training. The rightmost column illustrates the relative training time compared to standard training. The results highlight the effectiveness of our simple defense strategy, POD, in defending against previously unseen adversarial patches, with negligible additional training time. Notably, POD enhances detection precision in clean scenarios.
@@ 5.1 Evaluation for Digital and Physical Adversarial Patches
Additionally, despite POD not being exposed to AdvPatch during training, it achieves a comparable Average Precision against AdvPatch as Adv-POD, a model explicitly trained with AdvPatch. This may be attributed to Adv-POD suffering from overfitting and failing to generalize to the test samples. It is well-known that typical adversarial training suffers from adversarial overfitting [rice2020overfitting], wherein the model overfits to worst-case adversarial images within the training samples. In contrast, POD avoids overfitting due to its simplicity and randomness of augmentations, resulting in generalization against unseen strong attacks.
@@
[…]
POD improved precision in the clean scenario. Notably, the POD variants did not sacrifice detection precision in the clean scenario but rather improved it, in contrast to conventional adversarial training schemes [2, madry2017towards] that often suffer from a trade-off between clean and robust accuracy [zhang2019theoretically_trade_off]. This is attributed to the POD variants simulating various patch-based occlusions with diverse patch types with random size and location. We presume that the POD variants learn diverse features that are useful to robustly detect humans even when part of the body is occluded, and consequently, they improve accuracy in the clean scenario.

### v3

@@ Abstract
Infrared detection is an emerging technique for safety-critical tasks owing to its remarkable anti-interference capability. However, recent studies have revealed that it is vulnerable to physically-realizable adversarial patches, posing risks in its real-world applications. To address this problem, we are the first to investigate defense strategies against adversarial patch attacks on infrared detection, especially human detection. We propose a straightforward defense strategy, patch-based occlusion-aware detection (POD), which efficiently augments training samples with random patches and subsequently detects them. POD not only robustly detects people but also identifies adversarial patch locations. Surprisingly, while being extremely computationally efficient, POD easily generalizes to state-of-the-art adversarial patch attacks that are unseen during training. Furthermore, POD improves detection precision even in a clean (i.e., no-attack) situation due to the data augmentation effect. Our evaluation demonstrates that POD is robust to adversarial patches of various shapes and sizes. The effectiveness of our baseline approach is shown to be a viable defense mechanism for real-world infrared human detection systems, paving the way for exploring future research directions.
@@ 3.2 Detecting Patch-based Occlusions
Additionally, since the primary objective is to detect people robustly, we employ weighted cross-entropy loss [18] to prioritize the human class over the patch class:
@@ 4.4 Comparison between Adversarial Training Scheme
Furthermore, we developed a POD variant called “Adv-POD,” incorporating an adversarial training scheme that explicitly inputs adversarial patches during training, modified from Ad-YOLO [16]. Unlike Ad-YOLO, which is inadequate for defending against location-optimized patches like HCB and shape-loc due to its central patch placement, Adv-POD employs the same random patch placement scheme as POD. This modification allows us to fairly compare the effects of using basic patches (Sec. 3.1) or adversarial patches.
The procedure for Adv-POD comprises four steps: (1) a model is trained using a history of adversarial patches; (2) a new adversarial patch is generated every 15 epochs, starting from the $5_{th}$ epoch; (3) the adversarial patches are stored in the adversarial patch history for subsequent training; (4) the stored patches are randomly selected and attached to input image to train the detection model. We used the AdvPatch attack [9] for adversarial patch generation, positioning the patches in the exact same way as for POD, i.e., randomly determining their sizes and locations for fair comparison.
@@ 5 Results
In this section, we present experimental results demonstrating the effectiveness of POD against diverse adversarial patch attacks. The POD’s effectiveness is summarized in Table 3.
@@
Patch Type Patch Detect. Clean Noise Patch AdvPatch HCB Time (rel.)
@@
Table 1: Average precision (AP@0.5) for adversarial patches. Model names with “noDet” indicate models without the patch-detection module described in Sec. 3.2. Attacks marked with $\dagger$ are not physically realizable in the infrared scenario. Results marked with $\ast$ are the results for seen attacks: the adversarial patch was encountered during training. The rightmost column illustrates the relative training time compared to standard training. The results highlight the effectiveness of our simple defense strategy, POD, in defending against previously unseen adversarial patches, with negligible additional training time. Notably, POD enhances detection precision in clean scenarios.
@@ 5.1 Evaluation for Digital and Physical Adversarial Patches
Additionally, despite POD not being exposed to AdvPatch during training, it achieves a comparable Average Precision against AdvPatch as Adv-POD, a model explicitly trained with AdvPatch. This may be attributed to Adv-POD suffering from overfitting and failing to generalize to the test samples. It is well-known that typical adversarial training suffers from adversarial overfitting [23], wherein the model overfits to worst-case adversarial images within the training samples. In contrast, POD avoids overfitting due to its simplicity and randomness of augmentations, resulting in generalization against unseen strong attacks.
@@
POD achieved performance comparable to that of its adversarial variant, Adv-POD, yet with a notably shorter training time. The results in Table 1 demonstrate that POD and Adv-PODnoDet achieved similar performance in all scenarios, although the training time of POD is significantly shorter than that of Adv-PODnoDet. This is because, with the adversarial training scheme, it is time-consuming to generate adversarial examples during training, whereas POD relies solely on efficient augmentation techniques to simulate patch-based occlusions. Therefore, POD is an efficient and practical defense mechanism that can easily scale to train a model on a large dataset.
POD improved precision in the clean scenario. Notably, the POD variants did not sacrifice detection precision in the clean scenario but rather improved it, in contrast to conventional adversarial training schemes [2, 24] that often suffer from a trade-off between clean and robust accuracy [25]. This is attributed to the POD variants simulating various patch-based occlusions with diverse patch types with random size and location. We presume that the POD variants learn diverse features that are useful to robustly detect humans even when part of the body is occluded, and consequently, they improve accuracy in the clean scenario.